OpenStudy (anonymous):

two lines are perpendicular. if one line has a slope of 1/21 what is the slope of the other line

OpenStudy (solomonzelman):

if line `A` and `B` are perpendicular to each other, and line `A` has a slope of `m`, then line `B` would have a slope of `-1/m`

OpenStudy (anonymous):

so its -1/21

OpenStudy (solomonzelman):

not exactly

OpenStudy (solomonzelman):

in your case m=1/21 so -1/m = -1/(1/21)

OpenStudy (anonymous):

thx

OpenStudy (solomonzelman):

that is not it.

OpenStudy (solomonzelman):

you have to simplify that

OpenStudy (anonymous):

ok yea its 1/21

OpenStudy (anonymous):

or -21

OpenStudy (solomonzelman):